# QueStER: Query Specification for Generative Retrieval

### v1

@@ 4.2.1 Main results
In this section, we present the results of our model QueStER which is based on a Qwen-4B backbone, using cross-encoder assessments, a standard deviation $\nu=0.5$ when computing SoftNDCG and a cut-off at 10,000.666Note that because labeling with cross-encoder is costly, we assume that a document is not relevant if not in the top-100 re-labeled by the CE
@@
On MS MARCO Dev and TREC DL’19/’20 (Table 1), our GRPO-trained policy QueStER improves over BM25 and Qwen3-4B zero-shot prompting. Compared to more closely related approaches (Q2D, HyDE, LameR, MuGI), QueStER has strong effectiveness while avoiding their generation overhead and instability. Finally, compared to dual encoders (ANCE, SPLADEv2, ColBERTv2), which have been trained on MS MARCO, QueStER – as the other query rewriting approaches – have a much lower nDCG@10 (on DL’19 and ’20, is 71.7 for SPLADEv2 vs 62 for QueStER) but a competitive R@1K (78.9 vs 81.8).
@@
On OOD datasets BEIR thakur2021beir , QueStER outperforms all the neural IR baselines (SPLADEv2, ColBERTv2) as reported in Table 2 – showing the interest of such approaches for OOD retrieval. It is also competitive with query rewriting models (e.g., MuGI/GPT4 and Query2Doc), while being more efficient (see Section 4.2.2). By contrast, prompt-heavy methods (HyDE, MuGI, and LameR) rely on large proprietary LLMs and stochastic decoding (temperature $\tau>0$), so their outputs are expensive to compute (several subsets are omitted due to the generation cost) and scores vary across runs. We instead decode with temperature $\tau=0$, producing deterministic, fully reproducible rewrites.
@@ 4.2.2 Efficiency Analysis
Figure 2 summarizes the efficiency-effectiveness trade-off of the different models (using one thread when retrieving with Lucene, on the 43 queries from DL’19). Leaving out the generation processing time, BM25 has the best efficiency (16.3 ms/query) but is less effective (nDCG@10 is 50.6). Our model QueStER has a strong balance (62.3 nDCG@10 with 28 ms/query). Our smaller variants lag in quality despite better efficiency. Prompt-expansion methods such as MuGI and LameR achieve a higher nDCG@10 but at the cost of efficiency (more than 100 ms/query). When taking into account generation processing time, QueStER is much more efficient than alternatives – even though it is difficult to compare processing time since MuGI relies on OpenAI APIs (this explains why we do not report any number), the sheer difference in the number of parameters and on the number of samples needed (several for MuGI, one for QueStER) implies a much better efficiency for QueStER.
@@ 4.3 Ablations
RaFE (mao2024rafe) uses supervised fine-tuning (SFT) to warm up the model. To investigate the effect of SFT in our case, we conduct experiments in which we randomly sample 50,000 queries (from the MS MARCO training dataset, with no overlap with the queries used to train the model with RL). We use a large model (Qwen3-32B) to generate rewritten queries, generating 20 reformulated queries, keeping the best one if improving retrieval performance. These couples (query, reformulated query) are used as a supervised dataset.
@@
Following MuGI zhang2024exploring, we examine whether reusing the original user query helps retrieval. More precisely, we tried to add it after training (inference), but this only provides a negligible gain (+0.1 in nDCG@10). Adding the query during both training and inference has a slight negative effect (-0.2 in nDCG@10). However, both strategies slightly improve the consistency of the performance across datasets, but more work is needed to understand exactly the effect of this modification.
@@ 5 Conclusion and Future Work
In this paper, we present QueStER, a generative retrieval approach that rewrites queries, which is trained with GRPO with a reward based on SoftRank taylor2008softrank and distillation. We show that QueStER has the best tradeoff between efficiency and effectiveness in its class of models (LLM-based query rewriting) and is competitive out-of the domain with neural IR models, even if it still lags behind when in-domain. This shows the potential of this type of approach – we plan to explore more structured (and thus expressive) query languages as well as other optimization techniques in the future.

### v2

@@ 4.2.1 Main results
In this section, we present the results of our model QueStER which is based on a Qwen-4B backbone, using cross-encoder assessments, a standard deviation $\nu=0.5$ when computing SoftNDCG and a cut-off at 10,000.777Note that because labeling with cross-encoder is costly, we assume that a document is not relevant if not in the top-100 re-labeled by the CE
@@
On MS MARCO Dev and TREC DL’19/’20 (Table 1), our GRPO-trained policy QueStER improves over BM25 and Qwen3-4B zero-shot prompting. Compared to more closely related approaches (Q2D, HyDE, LameR, MuGI), QueStER has strong effectiveness while avoiding their generation overhead and instability. Compared to the state-of-the-art generative retrieval model like SEAL and MINDER, our model also achieve a better effectiveness, which illustrate our proposed methods could perform better to large scale dataset. Finally, compared to dual encoders (ANCE, SPLADEv2, ColBERTv2), which have been trained on MS MARCO, QueStER – as the other query rewriting approaches – have a much lower nDCG@10 (on DL’19 and ’20, is 71.7 for SPLADEv2 vs 62 for QueStER) but a competitive R@1K (78.9 vs 81.8).
@@
On OOD datasets BEIR Thakur et al. (2021) , QueStER outperforms all the neural IR baselines (SPLADEv2, ColBERTv2) as reported in Table 2 – showing the interest of such approaches for OOD retrieval. It is also competitive with query rewriting models (e.g., MuGI/GPT4 and Query2Doc), while being more efficient (see Section 4.2.2). By contrast, prompt-heavy methods (HyDE, MuGI, and LameR) rely on large proprietary LLMs and stochastic decoding (temperature $\tau>0$), so their outputs are expensive to compute (several subsets are omitted due to the generation cost) and scores vary across runs. We instead decode with temperature $\tau=0$, producing deterministic, fully reproducible rewrites.
@@ 4.2.2 Efficiency Analysis
Figure 2 summarizes the efficiency-effectiveness trade-off of the different models (using one thread when retrieving with Lucene, on the 43 queries from DL’19). Leaving out the generation processing time, BM25 has the best efficiency (16.3 ms/query) but is less effective (nDCG@10 is 50.6). Our model QueStER has a strong balance (62.3 nDCG@10 with 28 ms/query). Our smaller variants lag in quality despite better efficiency. Prompt-expansion methods such as MuGI and LameR achieve a higher nDCG@10 but at the cost of efficiency (more than 100 ms/query).
Generation time
QueStER generates a single keyword specification capped at 64 tokens using a small open model ($\leq 4B$ parameters). In contrast, MuGI and LameR rely on large API-based LLMs (e.g., GPT-3.5/4, $\sim$170B) and can generate substantially longer expansions; for instance, MuGI may produce 4–5 rewrites of up to 512 tokens each per input query. As a result, end-to-end generation latency is not directly comparable across methods, and Figure 2 reports retrieval-time latency only.
@@ 4.3 Ablations
RaFE (Mao et al., 2024) uses supervised fine-tuning (SFT) to warm up the model. To investigate the effect of SFT in our case, we conduct experiments in which we randomly sample 50,000 queries (from the MS MARCO training dataset, with no overlap with the queries used to train the model with RL). We use a large model (Qwen3-32B) to generate rewritten queries, generating 20 reformulated queries, keeping the best one if improving retrieval performance. These couples (query, reformulated query) are used as a supervised dataset.
@@
Following MuGI Zhang et al. (2024a), we examine whether reusing the original user query helps retrieval. More precisely, we tried to add it after training (inference), but this only provides a negligible gain (+0.1 in nDCG@10). Adding the query during both training and inference has a slight negative effect (-0.2 in nDCG@10). However, both strategies slightly improve the consistency of the performance across datasets, but more work is needed to understand exactly the effect of this modification.
@@ 5 Conclusion and Future Work
In this paper, we present QueStER, a generative retrieval approach that rewrites queries, which is trained with GRPO with a reward based on SoftRank Taylor et al. (2008) and distillation. We show that QueStER has the best tradeoff between efficiency and effectiveness in its class of models (LLM-based query rewriting) and is competitive out-of the domain with neural IR models, even if it still lags behind when in-domain. This shows the potential of this type of approach – we plan to explore more structured (and thus expressive) query languages as well as other optimization techniques in the future.
@@
Acknowledgments
The authors acknowledge Air Liquide and ANR - FRANCE (French National Research Agency) for its financial support of the GUIDANCE project n°ANR-23-IAS1-0003. This work was granted access to the HPC resources of IDRIS & CINES (Project No. A19) under the allocation made by GENCI.